# The Chemistry and Physics of 199Hg Nuclear Spin Polarization Relaxation in Quantum Magnetometry Cells

### v1

@@ I.2 Oxygen
Armed with this knowledge, we took steps to remove oxygen from the nitrogen buffer gas by bubbling the gas through liquid amalgams of Mg-Hg or Al-Hg followed by a liquid nitrogen trap to remove Hg vapor.[30] After such purification, the Hg density became constant; however, the cell lifetimes would rapidly degrade with resonance light irradiation.
A further study of photosensitized reactions indicated that metastable 6 3P0 atoms, which are readily formed with collisions between 6 3P1 atoms and nitrogen molecules, due to resonance with the rotation states of N2. The 6 3P0 excited atoms still carry considerable energy, see Table 1, and can damage the wall coatings. The metastable lifetime has been estimated as seconds, so there is ample time for diffusion to the cell wall.
@@
Indeed, even bare fused silica is prone to damage by these excited atoms. In our earlier work with testing spatial isotropy, we discovered that fused silica optical pumping cells that were irradiated while at room temperature would suffer irreversible damage as manifested by a shorted spin lifetime. The damaging effects have been known since 1930 [32] but apparently not by us.
There was a vague notion that hydrogen was responsible for the degradation of the spin lifetime; however, a model was never developed, and the notion tended to be dismissed because of the successful operation of hydrogen masers (more on this later).
The best option to quench 6 3P1 (that is, perturb the excited state causing a radiative transition or energy transfer) directly to the ground state and also to quench 6 3P0 (which is not quenched by nitrogen) is carbon monoxide, CO. Cells were made with 5% CO and 95%N2 and the cell life expectancy improved. UW 199Hg EDM experiments over the next decade employed such cells.
@@ I.3 Further Advances
The latest 199Hg EDM experiment is reported in [17], and incorporated a number of improvements. First, it used a pulsed pump and probe method that gave a factor of nearly 2 in the average atomic polarization, as suggested in [33]. This was demonstrated together with spin precession detected by the Faraday rotation of the off-resonant light provided by an 200Hg discharge lamp. (It should be noted that a similar system is described in [4].)
With the advent of 254 nm tunable lasers, a high-performance system was constructed, and the tightly focuses beam allowed the use of higher quantum efficiency small area photodiodes. The small diameter laser beams also meant that the cell walls were not fully bathed in resonance light, plus exposure to Hg was reduced by use of non-resonant detection light. A higher 199Hg density was also possible when using the Faraday effect and lasers, which could be tuned away from the line center.
The cells contained pure CO as a buffer gas, instead of a mixture with N2. Obviously, care was taken to remove oxygen; however, the level of hydrogen contamination is not reported.
These cells were carefully constructed with polished, optically flat end disks and cylinder edges, where a vacuum adhesive was applied. Instead of Varian Torr-Seal epoxy, Space Environment Labs Vacseal silicone-resin-based leak sealant was used. This material has a much lower outgassing rate compared to Torr-Seal, which, together with the close fit between the cylinder and the end-cap disks, minimized the exposure of contaminants that might be generated by the sealant. Furthermore, the cell components were carefully cleaned by refluxing in hydrochloric acid. These cells had stable lifetimes of at least 100 s. There is some indication that the cured silicone resin is slightly permeable to oxygen, which could be a reason for the cell lifetimes to be stabilized. [34]
@@ I.4 Hydrogen
Since the time that CO was identified as a useful buffer gas, there has been a serious suspicion that cell degradation is due to hydrogen. However, this notion was often met with skepticism, with a counterexample being the Hydrogen Maser: If H atoms were a destructive agent, how could a maser ever operate? In fact, this atomic H is a problem for masers.
@@
In the same year, Berg, one of the coauthors of this work, published a single-author paper stating that excess hydrogen from water-based Teflon coatings and non-UHV vacuum components, rubber o-rings, etc., should be eliminated and had empirical evidence to support that notion.[36]
Later work shows that atomic H actually degrades Teflon coatings, as evidenced by the damage at the atomic beam focus point, on the surface opposite the entrance aperture of the maser bulb.[37]
We encountered similar problems with 199Hg when we used a water-based Teflon coating; coatings made from pure PTFE (polytetrafluoroethylene) in a volatile non-polar carrier give stable and reproducible nuclear spin polarization lifetimes; however, the lifetimes are limited to 15-20 s per cm mfp.
@@ I.4.1 Autocatalytic Hydrogen Production
This leads us to surmise that once hydrogen is present in the system, the rate of generation of more hydrogen actually increases, in the presence of exited-state Hg. That is because each H2 molecule split by an interaction with an Hg will more likely remove another H from a wall coating molecule and form a new H2 molecule. Thus, the degradation process accelerates, and ultimately, the cells, which initially can be rejuvenated by ritualistic treatments involving melting the wax, eventually a useful performance time is shortened to the point that the cells become useless.
Furthermore, atomic H itself is a free radical, and its presence in the cell along with the buffer gas will cause spin relaxation of 199Hg . The lifetime resulting from this will be estimated in the Discussion section.
@@ I.5 Surface States
A previous study, with its carefully acquired large data set, [38], has been very helpful in understanding the dynamics of spin relaxation 199Hg . Here we add to the analysis presented there, first in regard to Eq. (3) of [ 38]. This equation is true if there is a single bound state for an 199Hg atom on the cell surface. However, the number of bound surface states grows with mass,[39], a point that has rarely, if ever, been taken into account in previous similar work (see, e.g., [40]).
The number of states and their energies can be estimated using the WKB approximation, assuming the van der Waals potential (short-range Casimir-Polder force) between an 199Hg atom and a dielectric surface as
@@
where $C\approx-6\times 10^{-4}\ e$V nm3 for a heavy polarizable atom near a dielectric surface, with $z$ the distance from the surface in nm and $d_{0}\approx 0.14$ nm representing a sum of radii of the Hg atom and atoms bound in the fused silica surface. The coefficient can be calculated using the formalism in [ 41] with the dielectric function for fused SiO2 in [ 42], Fig. 4, where the calculation was done for Cs. The value for Hg is about 1/6 of that for Cs, due to the shorter wavelength of the allowed electric dipole moment transition of Hg, 180 nm compared to 895 nm and 854 nm (D1 and D2 transitions) for Cs, together with the reduced dielectric response of SiO2 at shorter wavelengths, leading to a reduction in the overlap integral.
@@
where we allow $\tau_{0}$ to depend on $n$, and $E_{n}$ are the bound state energies. However, since each level is likely to have a different sticking or residence time, a more profitable approach is to calculate the fraction of atoms on the surface by using the partition function.
However, there is no need to consider a sticking time at this point. The binding energies of the levels are occupied according to the Boltzmann distribution, so we can determine the average fraction of the total number of atoms distributed among the bound states. Assume that the atoms move unhindered while in the main cell volume (away from the wall) and at the wall there are bound states with energy $-E_{n}$ (for the following, we take $E_{n}=|E_{n}|$). While in a bound state, an atom can move freely parallel to the surface. Let the volume and area of, for simplicity, a spherical cell be $V$ and $A$, $\beta=1/k_{b}T$, and $p$ be the momentum of an atom. Using simple statistical theory, the probability $\epsilon_{n}$ that an atom spends time in the surface state $n$ is (assuming that this probability is much lower than 1 ),
@@
where $\lambda_{T}=\hbar(2\pi\beta/m)^{1/2}$ is the thermal wavelength, $\beta=1/k_{b}T$, $m$ is the mass of an atom, and $R$ the radius of the cell, about 1 cm. For Hg at $293\mathrm{~{}K}$, $\lambda_{T}=$ $7\times 10^{-10}\mathrm{~{}cm}$ and using $E_{n}=0.21\ e$V with $R=1\mathrm{~{}cm}$ then $\epsilon=3\times 10^{-5}$.
@@
and the integral is the “exponential integral function," ${\rm Ei}(E_{max}/k_{b}T)-{\rm Ei}(E_{min}/k_{b}T)$, and using Eq. (5.1.10) of 43,
@@
The relaxation rate should depend on the time that an atom is stuck on the surface; however, it is unclear if that is the product of a sticking time and a Boltzmann factor. Given that the time between wall collisions is $\approx v/R$ which is proportional to $\sqrt{T}$, then the relaxation rate should be
@@
because $\lambda_{T}\propto 1/\sqrt{T}$. Noting that the derivative of the summation in our original integrand,
$\frac{1}{\Gamma_{p}}\frac{\partial\Gamma_{p}}{\partial T}=\frac{1}{T}-\frac{{e^{E_{max}/k_{b}T}}}{T}\left[{\sum_{j=1}^{\infty}\frac{(E_{max}/k_{b}T)^{j}}{j\cdot j!}}\right]^{-1}.$ (21)
From the graph Fig. 4 of [38], this ratio is about 0.03 at 260 K and 0.02 at 300 K, compared to the formula above numerically determined using $E_{max}=0.21$, with dielectric susceptibility the only specific material properties employed. Of course, correlations times, etc. will depend on temperature, but only weakly. Note that the collision rate times the sticking time $\Gamma_{p}\propto\sqrt{t}e^{\beta E_{0}}$ can be used to fit the data and approximately determines $E_{0}=0.15\ e$V which is close to $0.16\ e$V as obtained in [38] from the data presented in Fig. 4 of that paper.
@@
Note that we did not need to consider a sticking time at all. There is another time scale in the problem, which is the free flight time between wall collision, which for a 1 cm dimension cell is about $10^{-4}$ s. If each traverse across the cell results in a sticking collision, then the time at atom is stuck in the potential wells should be $10^{-4}\epsilon=6\times 10^{-9}$ s to maintain the statistically necessary Boltzmann distribution of atoms. The correlation times in [38] are slightly shorter than this, which we might expect since the atoms move along the surface, the only requirement is that the sticking time be longer than the correlation time; the data in [38] implies the opposite, that the correlation time was less than the sticking time and was left as an open question. If the probability of a wall collision has a probability less than one for an atom to stick, the dwell time becomes longer, as required by microscopic reversibility and the value of $\epsilon$. This implies that the approximately 6 ns time obtained above is strictly a lower limit.
@@
The spatial extent of the surface binding region is only about 0.5 nm, so in $6\times 10^{-9}$ s a 199Hg atom will suffer $10^{4}$ collisions with the wall. Given that there are about 100 surface states, a random walk out of the miasma could very well require $\sim 100^{2}$ steps.
@@
The long wall sticking time for 129Xe as reported in [40] can be explained by there being only a small probability of capture in the wall potential for each wall collision, unlike the case of 199Hg discussed above, where it seems to be of order 50%. Given the chemical reactivity of Hg compared to Xe, this is not surprising, so a sticking probability for Xe between 0.1 and 1 % is needed to explain the results in [40]. The reported binding energy of 0.1 $e$V as determined from the variation of the relaxation rate temperature is probably greater, as in the case of 199Hg where, without the full theory, 0.16 eV was obtained in [38] compared to $\approx 0.21$ eV for the most tightly bound surface obtained from the van der Waals interaction together with the WKB approximation.
@@
The idea of using the vibrational period of an atom near the wall as a sort of sticking time might be applicable for atoms with a bound state spectrum levels below $k_{b}T=0.025\ e$V. This idea was used in [44] from [45], and appears to not obviously apply to 199Hg or 129Xe.

### v2

@@ I.2 Oxygen
Armed with this knowledge, we took steps to remove oxygen from the nitrogen buffer gas by bubbling the gas through liquid amalgams of Mg-Hg or Al-Hg followed by a liquid nitrogen trap to remove Hg vapor.[39] After such purification, the Hg density became constant; however, the cell’s $\tau_{w}$ rapidly degraded with resonance light irradiation.
A further literature study of photosensitized reactions indicated that metastable 6 3P0 Hg atoms can be readily formed with collisions between 6 3P1 Hg atoms and nitrogen molecules, due to resonance with the rotational states of N2. The 6 3P0 excited atoms still carry considerable energy, see Table 1, and can damage the wall coatings. The metastable atomic state lifetime has been estimated as a few seconds, so there is ample time for diffusion to the cell wall.
@@
Indeed, even bare fused silica is prone to damage by these excited atoms. In our earlier work with testing spatial isotropy, we discovered that fused silica optical pumping cells that were irradiated while at room temperature would suffer irreversible damage as manifested by a shorted $\tau_{w}$. The damaging effects on silica have been known since 1930 [41] but apparently not by us.
There was a vague notion that hydrogen was responsible for the degradation of the $\tau_{w}$; however, a model was never developed and the notion tended to be dismissed due to the successful operation of hydrogen masers (more on this later).
The best option to quench 6 3P1 (that is, perturb the excited state causing a radiative transition or energy transfer) directly to the ground state and also to quench 6 3P0 (which is not quenched by nitrogen) is carbon monoxide, CO. Cells were made with 5% CO and 95% N2 and the useful life expectancy of cells under UV irratiation improved. The University of Washington 199Hg EDM experiments over the next decade employed such cells.
@@ I.3 Further Advances
The latest 199Hg EDM experiment is reported in [28], and incorporated several improvements. First, it used a pulsed pump and probe method that gave a factor of nearly 2 in the average atomic polarization. The spin precession was detected by use of the Faraday effect (linear polarization rotation) of second light beam that was not resonant with any 199Hg 254 nm transitions. Thus, there was little or no optical pumping relaxation due to the detection light, and the pump light was blocked during a measurement. (It should be noted that a similar system is described in [9].)
This high performance system was made possible by advent of 254 nm tunable lasers, and the tightly focused beams allowed the use of higher quantum efficiency small area photodiodes. The small diameter laser beams also meant that the cell walls were not fully bathed in resonance light, and exposure to Hg was reduced by the use of non-resonant detection light. A higher 199Hg density was also possible when using the Faraday effect and lasers, which could be tuned away from the line center. The cells contained pure CO as a buffer gas, instead of a mixture with N2. Obviously, care was taken to remove oxygen; however, the level of hydrogen contamination is not reported.
These cells were carefully constructed with polished, optically flat end disks and cylinder edges, where a vacuum adhesive was applied. Instead of Varian Torr-Seal epoxy, Space Environment Labs Vacseal silicone-resin-based leak sealant was used. This material has a much lower outgassing rate compared to Torr-Seal, which, together with the close fit between the cylinder and the end-cap disks, minimized the exposure of contaminants that might be generated by the sealant. The cell components were carefully cleaned by refluxing in hydrochloric acid. These cells had a stable $\tau_{w}$ of at least 100 s. This is notably lower than the 300 s light-limited spin polarization relaxation time obtained in the first EDM cells which had a small amount of O2 gas. Although the presence of O2 resulted in a loss in Hg (gas), it apparently improved the spin relaxation properties of the surface. It is well known that some cured silicone resins are slightly permeable to oxygen, [42] which could be a reason for the stabilization of the cell spin polarization relaxation time. [43]
@@ I.4 Hydrogen
Since CO was identified as a useful buffer gas, there has been a serious suspicion that the cell degradation is due to hydrogen. However, this notion was often met with skepticism, with a counterexample being the Hydrogen Maser: If H atoms were a destructive agent, how could a maser ever operate? In fact, this atomic H is a problem for masers.
@@
In the same year, Berg, one of the coauthors of this work, published a single-author follow-up paper stating that excess hydrogen from water-based Teflon coatings and non-UHV compatible vacuum components (rubber o rings, etc.) should be eliminated and had empirical evidence to support that notion.[45]
Later work shows that atomic H actually visibly degrades Teflon coatings, as evidenced by the damage at the atomic beam focus point, on the surface opposite the entrance aperture of the maser bulb.[46]
We encountered similar problems with 199Hg when we used a water-based Teflon coating; coatings made from pure PTFE (polytetrafluoroethylene) in a volatile non-polar carrier give stable and reproducible $\tau_{w}$; however, the $\tau_{w}$ are limited to 15-20 s per cm mfp.
@@ I.4.1 Autocatalytic Hydrogen Production
This leads us to surmise that once hydrogen is present in the system, the rate of generation of more hydrogen actually increases, in the presence of exited-state Hg. This is because the two H atoms from the splitting of each molecule split by an interaction with an Hg will more likely combine with and remove H atoms from a wall coating molecule and form new H2 molecules, rather than finding and recombining with their former partners. Thus, the degradation process accelerates and, ultimately, cells that initially could be rejuvenated by ritualistic treatments that involve melting wax, eventually had $\tau_{w}$ degraded to the point that the cells become useless.
Furthermore, atomic H itself is a free radical, and its presence in the cell along with the buffer gas can cause the relaxation of spin polarization of 199Hg , which adds to that due to the walls. The spin polarization relaxation time resulting from this is estimated in the discussion section.
@@ I.5 Surface States
A previous study, with its carefully acquired large data set, [47], has been very helpful in understanding the dynamics of spin relaxation of 199Hg . Here we add to the analysis presented there and recapitulate and expand the analysis presented in [48].
First, consider Eq. (3) of  [47]. This equation is true if there is a single bound state for an 199Hg atom on the cell surface. However, the number of bound surface states grows with the mass of the bound atom,[49], a point that has rarely, if ever, been taken into account in previous similar work (see, e.g., [50]).
The number of states and their energies can be estimated using the Wentzel-Kramers-Brillouin (WKB) approximation, which is an approximate method for determining the bound state energies and the wavefunctions of a massive particle bound in a spatially varying potential[51]. The van der Waals potential (short-range Casimir-Polder force) between an 199Hg atom and a dielectric surface is given by
@@
where $C\approx-6\times 10^{-4}\ e$V nm3 for a heavy polarizable atom near a dielectric surface, with $z$ the distance from the surface in nm and $d_{0}\approx 0.14$ nm representing the sum of radii of the Hg atom and atoms bound in the fused silica surface. The coefficient can be calculated using the formalism in  [52] with the dielectric function for fused SiO2 in  [53], Fig. 4, where the calculation was done for Cs. The value for Hg is about 1/6 of that for Cs, due to the shorter wavelength of the allowed electric dipole moment transition of Hg, 180 nm compared to 895 nm and 854 nm (D1 and D2 transitions) for Cs, together with the reduced dielectric response of SiO2 at shorter wavelengths. This leads to a reduction in the overlap integral.
@@
where we allow $\tau_{0}$ to depend on $n$, and $E_{n}$ are the bound state energies determined by the WKB approximation. However, since each level is likely to have a different sticking or residence time, a more profitable approach is to calculate the fraction of atoms on the surface by using the partition function. The binding energies of the levels are occupied according to the Boltzmann distribution, so we can determine the average fraction of the total number of atoms distributed among the bound states and scale that according to the time between wall collisions.
Assume that the atoms move unhindered while in the main cell volume (away from the wall) and at the wall there are bound states with energy $-E_{n}$ (for the following, we take $E_{n}=|E_{n}|$). In a bound-surface state, we can assume that an atom moves more or less freely parallel to the surface. The model generally assumed for this problem is that the atom hops between lightly bound regions.[3] For simplicity, assume a spherical cell with volume and area $V$ and $A$, $\beta=1/k_{b}T$, and $p$ is the momentum of the atom. Using simple statistical theory, the probability $\epsilon_{n}$ that an atom spends time in the surface state $n$ is (assuming that this probability is dynamic and very small),
@@
where $\lambda_{T}=\hbar(2\pi\beta/m)^{1/2}$ is the thermal wavelength, $\beta=1/k_{b}T$, $m$ is the mass of an atom, and $R$ the radius of the cell, about 1 cm. Approximately, for Hg at $293\mathrm{~{}K}$, $\lambda_{T}=$ $7\times 10^{-10}\mathrm{~{}cm}$. For the most strongly bound WKB state, $E_{n}\approx 0.21\ e$V, and with $R=1\mathrm{~{}cm}$ determines $\epsilon\approx 3\times 10^{-5}$.
@@
and the integral is the “exponential integral function," ${\rm Ei}(E_{max}/k_{b}T)-{\rm Ei}(E_{min}/k_{b}T)$, and using Eq. (5.1.10) of [54],
@@
The spin polarization relaxation rate should depend on the time that an atom resides on the surface; however, it is not clear whether $\tau_{w}$ is the product of a sticking time and a Boltzmann factor as has often been assumed. [3, 47] Given that the time between wall collisions is $\approx v/R$, which is proportional to $\sqrt{T}$, the relaxation rate should be
@@
because $\lambda_{T}\propto 1/\sqrt{T}$. Noting that the fractional change in relaxation rate is readily found by taking the derivative of Eq. (12),
$\frac{1}{\Gamma_{w}}\frac{\partial\Gamma_{w}}{\partial T}=\frac{1}{T}-\frac{{e^{E_{max}/k_{b}T}}}{T}\left[{\sum_{j=1}^{\infty}\frac{(E_{max}/k_{b}T)^{j}}{j\cdot j!}}\right]^{-1},$ (14)
and therefore a direct comparison can be made with the data presented in Fig. 4 of [47]. The experimental values in [47] are 0.03 at 260 K and 0.02 at 300 K, compared to the numerically determined values 0.033 and 0.024 from Eq. (14), respectively, using $E_{max}=0.21$ from the WKB approximation. This agreement is quite good, determined by the use of only the frequency-dependent dielectric properties of fused silica together with other measured cell parameters (temperature, distances); in some sense the calculation is ab initio. Of course, correlations times will depend on temperature but only as the velocity.
As an aside, it is of course possible to use the wall collision rate times the exponentially temperature dependent “sticking time," with $\Gamma_{w}\propto\sqrt{T}\tau_{0}e^{\beta E_{0}}$ in a fit to the data and approximately determine the “surface binding energy." The result is $E_{0}=0.15\ e$V which is close to $0.16\ e$V as obtained in [47] from the same data presented in Fig. 4 of that article. Then, to get the sticking time, a value of $\tau_{0}$ is required. In [47], this is somewhat arbitrarily chosen as the inverse vibrational frequency of the CH4-Hg van der Waals complex, $\tau_{0}=10^{-12}$ s [55]. The predicted sticking time at 293 K based on [47] is therefore $0.510^{-12}\ {\rm s}\ \times e^{0.16/.025}=$ 0.5 ns, which is approaching an order of magnitude lower than the observed correlation times of 2 ns.
Alternatively, it is also possible to use the WKB results to estimate this frequency, by analogy with a harmonic oscillator and taking the difference in energy of the two most strongly bound vibrational states. This is about 0.02 eV, corresponding to a period $2\times 10^{-13}$ s. This could be expected because the molecular cluster vibrational frequency is determined by the reduced mass, which is essentially the CH4 mass; hence, the WKB spectrum has a lower energy by $\sqrt{16/199}=0.28$, which brings the frequencies into agreement. Using the WKB results in an even larger problem, a sticking time of 0.06 ns. It makes no sense for the correlation time to be less than the sticking time, so this approach does not really work.
@@
However, we do not need to consider a sticking time at all, but a dwell time. There is another time scale in the problem, which is the free flight time between wall collision, and for a 1 cm dimension cell, it is about $10^{-4}$ s. If each traverse across the cell results in a “sticking" collision, then the time an atom dwells in the potential well should be $10^{-4}\ {\rm s}\times\epsilon=6\times 10^{-9}$ s to maintain the statistically necessary Boltzmann distribution of atoms. The correlation times in [47] are slightly shorter than this, which might be expected because the atoms move along the surface while in a bound state and possibly react coherently only briefly with a given relaxing site. The only requirement is that the dwell time be longer than the correlation time.
The data as interpreted in [47] implies the opposite, that the correlation time is longer than the sticking time, and this impossibility was left as an open question. If the probability that a wall collision results in an atom sticking is very small, the dwell time necessarily becomes longer, as required by microscopic reversibility and the value of $\epsilon$. This implies that the approximately 6 ns dwell time obtained above is strictly a lower limit.
@@
The spatial extent of the surface binding region is only about 0.5 nm, so in $6\times 10^{-9}$ s, while oscillating in the potential well, an 199Hg atom will suffer $10^{4}$ collisions with the wall. Given that there are about 100 surface states, a random walk out of the miasma of surface states could very well require $\sim 100^{2}$ steps, with $100^{2}\times 6\times 10^{-9}$ s $=6\times 10^{-5}$ s, which can be used to estimate the probability of wall collision sticking. For $10^{-4}$ s free flight time between wall collisions, the probability is $0.66\times 10^{-5}\ {\rm s}/10^{-4}\ {\rm s}=$ 0.6.
@@
The long wall sticking time for 129Xe as reported in [50] can be explained if there is only a small probability of capture in the wall potential for each wall collision, unlike the case of 199Hg discussed above, where it is found to be of order 60%. Given the chemical reactivity of Hg in comparison to Xe, this is not surprising, so the sticking probability for Xe is likely smaller, and the interaction with relaxing sites might be reduced. Both effects are needed to explain the results in [50]. The reported binding energy of 0.1 $e$V, determined from the variation of the relaxation rate as a function of temperature, is probably too small. In the case of 199Hg where, without the full theory, 0.16 eV was obtained in [47] compared to $\approx 0.21$ eV for the most tightly bound surface obtained from the van der Waals interaction together with the WKB approximation, and the same conclusions about a “true" larger binding energy can be drawn in the case of Xe.
@@
The possibility of using the vibrational period of an atom near the wall as a sort of sticking time is not without merit. For example, the vibrational period was used in the analysis of the relaxation of the electron spin polarization of alkali atoms as described in [56], based on the surface physics considerations discussed in [57].